OpenStudy (anonymous):

simple cuts is a hair salon that offers two types of haircuts: a buzz cut and a flat-top. Jerry charges $9 for a buzz cut and $13 for a flat-top. If Jerry made $600 one day and knows that he three times as many buzzes as flat-tops, how many flat-tops did he cut that day? A. 15 B.21 C.40 D.280

OpenStudy (stamp):

define buzzes b, flat-tops f. b = 3f 9b + 13f = 600 9(3f) + 13f = 600 /* substitute b */ 27f + 13f = 600 /* simplify expression */ 40f = 600 /* next solve for f */